35=1/3b(5) solve for b v=1/3bh
Answer by nerdybill(7003)

35=1/3b(5) solve for b v=1/3bh
.
35=1/3b(5)
dividing both sides by 5:
7=1/3b
multiplying both sides by 3:
21 = b